OpenStudy (anonymous):

Given the functions f(n) = 25 and g(n) = 3(n − 1), combine them to create an arithmetic sequence, an, and solve for the 12th term.

OpenStudy (anonymous):

an = 25 − 3(n − 1); a12 = −11 an = 25 − 3(n − 1); a12 = −8 an = 25 + 3(n − 1); a12 = 58 an = 25 + 3(n − 1); a12 = 61

OpenStudy (anonymous):

wow does "combine" mean "add"?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

tell your math teacher there is no such mathematical operation as "combine"

OpenStudy (anonymous):

Lol, will do!!

OpenStudy (anonymous):

although you can "combine like terms" for example

OpenStudy (anonymous):

if it means "add' then stick a plus sign between them that is all

OpenStudy (anonymous):

Ok, so now have C and D left

OpenStudy (anonymous):

then replace \(n\) by \(12\) and see what you get

OpenStudy (anonymous):

58! So, is it C?